RC Meeting: August 28th, 2020

From clas12-run
Jump to: navigation, search

Back to Run Group F | Daily Run Group F Coordination Meetings


2020

March
19, 20, 21, 22, 23, 24
July
31
August
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31
September
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21

RC: M. Hattawy

Summary of Aug. 28th

Owl shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3836017
  2. Issues: lots of trips and complaints about the RFs. beam is down from 5 AM to work the south linac ioc
  3. Recorded data: 16M on D2

Day shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3836231
  2. Issues: beam is down because of the SL
  3. Studies: cleaned the beam pipe between the end of the CVT carte and 2H01A.
  4. Recorded data: 0

Swing shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3836427
  2. Studies: beam came back at 22:00 pm for 30 minutes. tuned the beam on the tagger, then on FC.
  3. Recorded data: 0

Whole day summary

  1. Total recorded data today: 16 M
  2. Part-II total accumulated data on D2: 1156M.
  3. Concerns: no new concerns, still we dont now what is the status of the rates, still scanning the target before going to production conditions.


Tomorrow daily RC meeting will be at usual time 12:00 PM using the link https://bluejeans.com/7576835804/

Recorded data numbers

Number of events recorded